In 1921, President Taft decided that the ratio of the length to the height of an American flag should be 1 to 1.9. If a flag is 25 square feet, how long and tall is it?
Answer by edjones(8007)

L*W=A
x*1.9x=25
1.9x^2=25
x^2=25/1.9=13.1579
x=3.6274' tall
1.9*x=6.8920' long
Ed